OpenStudy (anonymous):

On Friday it snowed 3 4/7 inches. On Saturday it snowed 2/7 inches. How much more did it snow on Friday than on Saturday

OpenStudy (anonymous):

In order to solve this we need to find the difference between the two numbers, i.e. we need to subtract: \[3\frac{4}{7} - \frac{2}{7} = ?\] Lucky for us both fractions already have a common denominator so we can just subtract them: \[3\frac{4}{7} - \frac{2}{7} = ?\]\[\frac{25}{7} - \frac{2}{7} = ?\]\[\frac{23}{7}\] Or in simplest terms: \[3\frac{2}{7}\]

OpenStudy (anonymous):

thx